OpenStudy (anonymous):

Which proportion has a solution of 8? 10/15 = n/12 10/15/=12/n 10/n =12/15 10/12 = n/15

OpenStudy (anonymous):

10/15=n/12 Because on further simplifying the equation and solving for n, we have 2/3=n/12 => n=24/3 => n=8